OpenStudy (anonymous):

Solve each system of equations. 1. x-2y+z=8 y-z=4 z=3 I've tried this problem and 2 other problems maybe 10 times each and I just can't figure it out. I don't know what I'm doing wrong. help.

OpenStudy (anonymous):

z=3 use this in equation #2 y-3=4 y=7

OpenStudy (anonymous):

z=3,y=7 use these in equation #1 x -2(7)+3=8 x-14+3=8 x-11=8 x=19

OpenStudy (anonymous):

z=3,y=7,x=19

OpenStudy (texaschic101):

check... x - 2y + z = 8 19 - 2(7) + 3 = 8 19 - 14 + 3 = 8 5 + 3 = 8 8 = 8 (correct) x = 19, y = 7, z = 3

OpenStudy (anonymous):

Thank you @d3v13 and @texaschic101 can you help me on 2 others?

OpenStudy (texaschic101):

what ya got

OpenStudy (anonymous):

2x-y+z=15 -6x-y-z=-11 4x-3y-z=0

OpenStudy (anonymous):

@texaschic101

OpenStudy (kitkat16):

z=11 x=1/2 y=-3